OpenStudy (anonymous):

a number is increased 80% or 480 what is the original number

OpenStudy (dan815):

480=0.8*x

OpenStudy (dan815):

what must the original number be

OpenStudy (goformit100):

Question is typed Wrong.... check it Ma'am

OpenStudy (mathstudent55):

Are you asking: 480 is 80% of a number. What is the number?

OpenStudy (goformit100):

it would be of in place of or...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

jeez the question says "a number is increased 80% or 480 what is the original number"

OpenStudy (anonymous):

translate as \(480\) is \(80\%\) of what number? put in math as \[480=.8x\] making \[x=480\div .8\]

OpenStudy (anonymous):

i am not really smart can you explain diffrent or give answer for meadle @satellite73

OpenStudy (mathstudent55):

@satellite73 The question is easier to understand now, when you saw it, because it has been edited. The original question had some words missing and was not as easy to understand.

OpenStudy (anonymous):

oh ok as for being really smart this is math and has nothing to do with being smart or not, it is just math it is a matter of interpreting what the question is actually asking it is worded funny

OpenStudy (anonymous):

i would rewrite a number is increased 80% or 480 what is the original number as a number is increased by 80% it was increased by 480 what was the original number

OpenStudy (anonymous):

which we could further translate as 480 is 80% of what?

OpenStudy (anonymous):

so how do i figure it out

OpenStudy (anonymous):

we don't know the original number (lets call it \(x\) ) but we know that \(80\%\) of it is \(480\) so \[.8\times x=480\]

OpenStudy (anonymous):

because \(80\%\) means \(\frac{80}{100}=.8\) and "of" means "times"

OpenStudy (anonymous):

since \[.8x=480\] which can find \(x\) by dividing \(480\) by \(.8\) which, if you like, you can do with a calculator

OpenStudy (anonymous):

\[480\div .8==4800\div 8=600\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw hope it is somewhat more clear

OpenStudy (anonymous):

yes